OpenStudy (anonymous):

You roll a normal 6-sided die. What is the probability that you will roll a 1 or a 4 or a 6?

Directrix (directrix):

If you toss a single die, there are six outcomes: 1, 2, 3, 4, 5, and 6. Probability is the number of successful outcomes divided by the number of possible outcomes. 1,4, and 6 are three of the possible successful (successful is what you want to happen) outcomes out of the six possible outcomes.

OpenStudy (ikram002p):

its simly the sum(sum cuz u have OR) of the the possible successful of ech one p(1,4,6)=p(1)+p(4)+p(6) =1/6+1/6+1/6 (since its a normal die =3/6= 0.5